OpenStudy (anonymous):

Verify the trig equations match by substituing identities. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^(2) x

OpenStudy (anonymous):

\[(\sin x)([\sin x/\cos x] * \cos x - [\cos x/\sin x] * \cos x) = 1-2\cos ^{2}x\ That's as far as I got